A circuit consists of a 12.0-VV battery connected to three resistors (44 ΩΩ , 19 ΩΩ and 110 ΩΩ ) in series.

1. Find the current that flows through the battery.

2.Find the potential difference across the 44 Ω resistor.

3.Find the potential difference across the 19 Ω resistor.

4.Find the potential difference across the 110 Ω resistor.
Answer #1

A) TOTAL Resistance = 44+19+110 = 173 ohms

Current = V/R = 12 / 173 = 0.0693 A = 69.3 mA

B) V = I*R = 0.0693 * 44 = 3.0492 V

C) V = I*R = 0.0693 * 19 = 1.3167 V

D) V = I*R = 0.0693 * 110 = 7.6230 V